# Banzhida: Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training

### v1

@@ Abstract
Large language models have achieved remarkable progress across many languages. However, Tibetan, as a representative low-resource language, is particularly underrepresented in existing models due to the scarcity of high-quality training corpora. To address this gap, we curate the largest Tibetan pre-training corpus to date, aggregating data from diverse sources and applying a dedicated data cleaning and processing pipeline tailored for Tibetan. With the curated data, we continue pre/post-training a multilingual base model into Banzhida, a multilingual large language model that advances generative AI for Tibetan. To evaluate the Tibetan capabilities of the model, we create new high-quality Tibetan benchmarks, and complement them with existing public benchmarks. Experimental results demonstrate that Banzhida consistently and significantly outperforms both open-source models of similar scale and Tibetan-tailored models across a wide range of tasks.
Banzhida: Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training
Leiyu Pan1, Bojian Xiong1, Lei Yang1, Renren Jin1,3, Shaowei Zhang1 Yue Chen1, Ling Shi1, Jiang Zhou1, Junru Wu1, Zhen Wang1, Jianxiang Peng1 Juesi Xiao1, Tianyu Dong1, Zhuowen Han1, Zhuo Chen1, Sangjee Dondrub2, Caizang Tai2 Haixing Zhao2, Huaque Cairang2, Suonan Cairang2, Rou Te2, Lengben Zhaxi2, Gazang Zhaxi2 Zhonglin Ye2, Yuhui Zheng2, Chunyan Peng2, Secha Jia2, Pema Tashi2, Cizhen Jiacuo2 Pema Dorjee2, Hongkai Liu2, Pema Yanggon2, Tsehang Dorjee2, Jiaxin Han2, Qiongying Hu2 Jilin Man2, Huanke You2, Yuqi Ren1,3, Duo La2222Project Lead., Deyi Xiong1,3111Corresponding author. [^2] Project Lead. 1TJUNLP Lab, Tianjin University 2Qinghai Normal University, 3Yuanhui AI Lab
@@ 1 Introduction
Tibetan, as a representative low-resource language, is one of China’s minority languages and is spoken by millions of native speakers. It serves as a vital vehicle for the preservation of Tibetan culture, religious practices, and historical heritage Liang et al. (2024); Huang et al. (2025). Despite its significant population base and cultural value, Tibetan remains severely underrepresented in terms of digitized linguistic resources and has long been categorized as a low-resource language Liu et al. (2022); Liu and Best (2025). In recent years, the development of multilingual LLMs has led to gradual improvements in support for Tibetan Liu et al. (2024); Lv et al. (2025). However, due to the dominance of high-resource languages in training corpora and the limited scale, quality, and standardization of Tibetan data, existing models still struggle with tasks involving Tibetan understanding and generation Wenhao et al. (2024).
In light of the aforementioned challenges, we create Banzhida, a multilingual LLM supporting Tibetan, Chinese, and English. We present a comprehensive pipeline for building a Tibetan-capable LLM, covering data curation, pre-training, post-training, and evaluation.
@@
To build Banzhida, we continue to pre-train Qwen2.5-7B-base model Qwen et al. (2025), a strong open-source multilingual LLM, to extend its capabilities to Tibetan. To facilitate effective cross-lingual transfer and accelerate the model’s adaptation, we incorporate both Chinese and English data alongside the curated Tibetan data during this continual pre-training phase. Following pre-training, we perform instruction tuning using multilingual instruction-following datasets collected from open-source repositories in Tibetan, Chinese and English, further enhancing ability of Banzhida to handle diverse downstream tasks across the three languages.
@@
  • With the curated data, we build Banzhida, a multilingual LLM with strong capabilities in Tibetan.
@@ 3.1 Data Collection
Although the open-source data we have collected includes a wealth of web-based content, it suffers from limited timeliness, with many recent materials not reflected in existing datasets. To address this limitation, we utilize the crawl4ai tool [^1] https://github.com/unclecode/crawl4ai to collect data from 55 Tibetan-language websites, resulting in a total of 204,789 webpages. The crawled data spans a wide range of domains, which are primarily on news, but also include content from books, Wikipedia, and other categories, contributing to the diversity of our Tibetan corpus. Detailed information of web data crawling is given in Appendix C.
@@
In addition to publicly available data, we also utilize a set of private Tibetan-language digital resources, primarily consisting of books and religion materials. These resources are obtained from non-public but legally accessible sources, such as internal digital archives and institutional collaborations. All data used in this category contain no personally identifiable information or user-generated content, and are solely used for research and model training purposes.
@@ 6 Experiments
Due to the current lack of standardized Tibetan benchmarks for LLMs, we constructed four benchmarks to evaluate general Tibetan capabilities. Combined with two existing benchmarks for traditional Tibetan NLP tasks, this enables a comprehensive evaluation of model performance. Experimental results demonstrate that Banzhida outperforms both open-source models of similar scale and Tibetan-tailored models across all evaluated tasks.
@@ 6.2 Evaluation Benchmarks
Banzhida-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15
Banzhida-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52
Banzhida-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15
@@
Banzhida-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15 13.81 9.20
Banzhida-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52 20.86 13.56
Banzhida-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15 21.56 11.42
@@ 6.3 Comparison against Open-Source Models
We conducted a comparative evaluation of three versions of our model: Banzhida-Base, which is trained with one stage of continual pretraining; Banzhida-Base-32k, which undergoes two stages of continual pretraining; and Banzhida-SFT, which is further instruction-tuned. These models are evaluated against leading general-purpose open-source models on both standard reasoning and Tibetan-specific benchmarks.
@@ 6.4 Comparison against Existing Tibetan-Tailored Models
We further compared our models with several instruction-tuned models specifically designed for Tibetan. As shown in Table 3, Banzhida consistently outperform existing Tibetan instruction-tuned models across all tasks. In particular, Banzhida-SFT achieves the strongest overall performance, showing substantial improvements in reading comprehension and question answering, where other models fall significantly behind. Moreover, in both zh-bo and en-bo translation tasks, Banzhida models achieve the highest BLEU scores among all evaluated models, demonstrating strong multilingual alignment and effective cross-lingual generalization. It is worth noting that our comparison only includes open-source Tibetan-tailored models. Models such as Sun-Shine Huang et al. (2025), which are not publicly available, are not included in this evaluation.
@@ 7 Conclusion
To address the limited support for Tibetan in existing models, we construct a high-quality, large-scale Tibetan pre-training corpus. With the curated corpus, we perform continual pretraining to develop Banzhida, a multilingual language model supporting Tibetan, Chinese, and English. Banzhida demonstrates superior performance compared to both general-purpose open-source models and Tibetan-tailored models of similar scale on our constructed Tibetan benchmark as well as traditional Tibetan NLP tasks.
@@ Appendix B Open-Source Data Collection Details
The open-source Tibetan corpus we collected primarily comes from two sources: (1) the Tibetan portions of multilingual open-source datasets, and (2) dedicated monolingual Tibetan datasets. These corpora cover a wide range of domains. However, we observe that a portion of the web-based data originates from different snapshots of Common Crawl, which may introduce redundancy. To address this, we appliy deduplication to remove duplicate content across time. A detailed breakdown of the collected open-source Tibetan corpora is provided in Table 5.
@@ Appendix C Web Data Crawling Details
We manually survey 55 Tibetan-language websites and collect web content from them. Most of these sites are news-oriented, and our primary goal is to obtain up-to-date Tibetan text with strong temporal relevance. During the web crawling process, we extract hyperlink elements from each page and determine whether the linked page belonged to the same root domain. If so, and the page has not been previously visited, we recursively crawl the subpage. Throughout the entire process, all crawling behavior strictly adhere to the rules defined in each site’s robots.txt file. A detailed list of the crawled websites and the number of pages collected from each is provided in Table 6.
Name Domain Size Date
Community Oscar Brack et al. (2024) Web 14.95 G 2024-12
CulturaX Nguyen et al. (2024a) Web 709.43 M 2023-09
HPLT-v2 Burchell et al. (2025) Web 1.86 G 2025-03
MaLA Corpus Ji et al. (2024) Mutli 2.77 G 2024-09
MC2 Zhang et al. (2024) Web 1.55 G 2023-12
Tibetan News Zhang et al. (2022) News 1.60 G 2023-08
Buddhist Digital Resource Books 3.15 G 2020-01
Table 5: Overview of collected open-source Tibetan datasets.
Root Domain Count Root Domain Count
http://ti.kbcmw.com 17,949 http://zwpaper.kbcmw.com 130
http://www.tibetology.ac.cn 4,591 http://tb.chinatibetnews.com 2
http://www.himalayabon.com 5,008 http://epaper.tibet3.com 27
http://tibet.people.com.cn 751 http://www.kangbatv.com 206
http://www.tibetanecology.org 1222 https://mewe.com 53
http://res.kangbatv.com 117 http://www.teducn.com 237
http://www.tbmgar.com 11,172 https://mix.com 79
http://www.shangri-latibet.cn 9,403 http://tb.tibet.cn 18,675
http://www.amdotibet.cn 30,080 http://tb.kangbatv.com 950
https://adarshah.org 78 http://ti.tibet3.com 38
https://online.adarshah.org 124 http://www.tibetcnr.com 899
http://xizang.news.cn 453 http://www.vtibet.com 1
http://www.qhfzb.com 9 http://my.haibei.gov.cn 3,483
https://tb.xzxw.com 30,280 http://www.qinghai.gov.cn 139
http://epaper.chinatibetnews.com 129 https://epaper.tibet3.com 468
http://www.yongzin.com 18 https://ti.zangdiyg.com 120
http://www.qhtibetan.com 873 https://tibet.cpc.people.com.cn 1
https://www.tibetcm.com 21,106 https://tb.chinatibetnews.com 2
http://zw.guoluo.gov.cn 14,939 https://www.qhfzb.com 11
http://www.qhtb.cn 16,119 https://epaper.chinatibetnews.com 5
http://www.tibet3.cn 91 https://www.himalayabon.com 354
http://www.jmjzjy.com 7,140 http://jinrizangxiang.com 1,794
http://www.filmtbt.com 876 http://www.gnu.org 19
http://www.tibetitw.com 1 http://savannah.gnu.org 1
http://ti.zangdiyg.com 1,905 https://tb.kangbatv.com 1,129
http://www.tibet3.com 1 https://ti.tibet3.com 1,208
http://tibet.cpc.people.com.cn 3 http://www.qiongbuwang.com 306
http://www.abadaily.com 14
Table 6: List of crawled websites and the corresponding number of webpages collected.
@@ Appendix D Quality Filter Details
To eliminate politically sensitive content from the collected data, we manually curated a blacklist of sensitive terms, as shown in Table 1. Any document containing one or more terms from this list is filtered out.
Figure 1: List of sensitive Tibetan words and expressions.
@@ Appendix E Continual Pre-training Details
To improve the training efficiency of the Banzhida model, we adopt a packing strategy Raffel et al. (2020) to preprocess the data. This involves concatenating and truncating the data to ensure that each training sample is exactly 4096 tokens in length. Given the large volume of training data, we also apply a pre-tokenization strategy, where the tokenized and packed datasets are preprocessed and saved in advance to streamline the subsequent training process.
@@ Appendix F Long Context Pre-training Details
Data Distribution. To fully account for the distributional characteristics of the training data in the first stage and further enhance the model’s long context processing capabilities in Tibetan, we introduce longer mixed-language texts (Chinese, English, and Tibetan) as training corpora in the second training stage. As shown in Table 7, we sample 0.56B tokens of Chinese data, 2.5B tokens of English data, and 1.5B tokens of Tibetan data (unused in the first stage) from long context sources in each respective language, resulting in a trilingual long context training set totaling 4.56B tokens.
@@
Language Name Size Usage
Chinese LongData- Corpus Yu et al. (2023) 36.0 G 0.56 B
Chinese-Reasoning-Distil-Data Zhang (2025) 1.5 G
Chinese-DeepSeek-R1-Distill-data-110k Cong et al. (2025) 0.6 G
English LongCorpus- 2.5B Chen et al. (2023) 9.3 G 2.5 B
OpenThoughts2-1M Guha et al. (2025) 7.8 G
Tibetan - 9.6 G 1.5 B
Table 7: Long context pre-training data details for three languages.
@@ Appendix G Post Training Details
We extensively collect open-source instruction datasets in Chinese, English, and Tibetan, as summarized in Table 8. Upon analysis, we find that most existing Tibetan instruction datasets are generated by translating open-source Chinese and English instruction data using Google Translate. While this approach is simple and convenient, it often results in low-quality Tibetan instructions due to the current limitations of Google’s Tibetan translation capabilities. Additionally, such translations may introduce unintended biases. To address these issues, and in addition to the datasets listed in Table 8, we collaborate with native Tibetan speakers to produce additional high-quality translation and summarization data. These samples are organized using instruction-style templates to further enhance the model’s instruction-following abilities in Tibetan, particularly in the areas of translation and summarization.
@@
Name URL Language
GLM-4-Instruct-4K-zh huggingface.co/datasets/rqq/GLM-4-Instruct-4K-zh Chinese
Align-Anything-Instruction-100K-zh huggingface.co/datasets/PKU-Alignment/Align-Anything-Instruction-100K Chinese
stem_zh_instruction huggingface.co/datasets/hfl/stem_zh_instruction Chinese
alpaca-gpt4-data-zh huggingface.co/datasets/llm-wizard/alpaca-gpt4-data-zh Chinese
TaP Jin et al. (2025) - Chinese
alpaca-cleaned huggingface.co/datasets/yahma/alpaca-cleaned English
Infinity-Instruct huggingface.co/datasets/BAAI/Infinity-Instruct English
tulu-3-sft-personas-instruction-following huggingface.co/datasets/allenai/tulu-3-sft-personas-instruction-following English
ultrachat_200k huggingface.co/datasets/HuggingFaceH4/ultrachat_200k English
TibetanSft_corpus huggingface.co/datasets/shajiu/TibetanSft_corpus Tibetan
Table 8: Post-training open source instruction data for three languages.
@@ Appendix I Evaluation Benchmarks Details
Appendix K Generations from Banzhida
In this section, we present a collection of Banzhida generated question–answer examples to illustrate the multilingual capabilities of our Tibetan large language model.

### v3

# Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training

@@ Abstract
Large language models have achieved remarkable progress across many languages. However, Tibetan, as a representative low-resource language, is particularly underrepresented in existing models due to the scarcity of high-quality training corpora. To address this gap, we curate the largest Tibetan pre-training corpus to date, aggregating data from diverse sources and applying a dedicated data cleaning and processing pipeline tailored for Tibetan. With the curated data, we continue pre/post-training a multilingual base model to enhance its generative capabilities in Tibetan. To evaluate the Tibetan capabilities of the model, we create new high-quality Tibetan benchmarks, and complement them with existing public benchmarks. Experimental results demonstrate that our model consistently and significantly outperforms both open-source models of similar scale and Tibetan-tailored models across a wide range of tasks.
Advancing Large Language Models for Tibetan with Curated Data and Continual Pre-Training
Leiyu Pan1, Bojian Xiong1, Lei Yang1, Renren Jin1,2, Shaowei Zhang1 Yue Chen1, Ling Shi1, Jiang Zhou1, Junru Wu1, Zhen Wang1, Jianxiang Peng1 Juesi Xiao1, Tianyu Dong1, Zhuowen Han1, Zhuo Chen1, Yuqi Ren1,2, Deyi Xiong1,2111Corresponding author. [^2] Project Lead. 1TJUNLP Lab, Tianjin University 2Yuanhui AI Lab
@@ 1 Introduction
Tibetan, as a representative low-resource language, is one of China’s minority languages and is spoken by millions of native speakers. It serves as a vital vehicle for the preservation of Tibetan culture and historical heritage Liang et al. (2024); Huang et al. (2025). Despite its significant population base and cultural value, Tibetan remains severely underrepresented in terms of digitized linguistic resources and has long been categorized as a low-resource language Liu et al. (2022); Liu and Best (2025). In recent years, the development of multilingual LLMs has led to gradual improvements in support for Tibetan Liu et al. (2024); Lv et al. (2025). However, due to the dominance of high-resource languages in training corpora and the limited scale, quality, and standardization of Tibetan data, existing models still struggle with tasks involving Tibetan understanding and generation Wenhao et al. (2024).
In light of the aforementioned challenges, we create a multilingual LLM supporting Tibetan, Chinese, and English. We present a comprehensive pipeline for building a Tibetan-capable LLM, covering data curation, pre-training, post-training, and evaluation.
@@
To build our model, we continue to pre-train Qwen2.5-7B-base model Qwen et al. (2025), a strong open-source multilingual LLM, to extend its capabilities to Tibetan. To facilitate effective cross-lingual transfer and accelerate the model’s adaptation, we incorporate both Chinese and English data alongside the curated Tibetan data during this continual pre-training phase. Following pre-training, we perform instruction tuning using multilingual instruction-following datasets collected from open-source repositories in Tibetan, Chinese and English, further enhancing ability of the model to handle diverse downstream tasks across the three languages.
@@
  • With the curated data, we build a multilingual LLM with strong capabilities in Tibetan.
@@ 3.1 Data Collection
Although the open-source data we have collected includes a wealth of web-based content, it suffers from limited timeliness, with many recent materials not reflected in existing datasets. To address this limitation, we utilize the crawl4ai tool [^1] https://github.com/unclecode/crawl4ai to collect data from 55 Tibetan-language websites, resulting in a total of 204,125 webpages. The crawled data spans a wide range of domains, which are primarily on news, but also include content from books, Wikipedia, and other categories, contributing to the diversity of our Tibetan corpus. Detailed information of web data crawling is given in Appendix C.
@@
In addition to publicly available data, we also utilize a set of private Tibetan-language digital resources, primarily consisting of books and classical materials. These resources are obtained from non-public but legally accessible sources, such as internal digital archives and institutional collaborations. All data used in this category contain no personally identifiable information or user-generated content, and are solely used for research and model training purposes.
@@ 6 Experiments
Due to the current lack of standardized Tibetan benchmarks for LLMs, we constructed four benchmarks to evaluate general Tibetan capabilities. Combined with two existing benchmarks for traditional Tibetan NLP tasks, this enables a comprehensive evaluation of model performance. Experimental results demonstrate that our model outperforms both open-source models of similar scale and Tibetan-tailored models across all evaluated tasks.
@@ 6.2 Evaluation Benchmarks
Ours-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15
Ours-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52
Ours-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15
@@
Ours-Base 30.30 36.20 44.32 59.80 60.80 49.43 66.15 13.81 9.20
Ours-Base-32k 30.09 36.58 44.62 58.60 60.60 38.42 55.52 20.86 13.56
Ours-SFT 30.64 36.71 48.39 57.80 61.86 46.15 63.15 21.56 11.42
@@ 6.3 Comparison against Open-Source Models
We conducted a comparative evaluation of three versions of our model: Ours-Base, which is trained with one stage of continual pretraining; Ours-Base-32k, which undergoes two stages of continual pretraining; and Ours-SFT, which is further instruction-tuned. These models are evaluated against leading general-purpose open-source models on both standard reasoning and Tibetan-specific benchmarks.
@@ 6.4 Comparison against Existing Tibetan-Tailored Models
We further compared our models with several instruction-tuned models specifically designed for Tibetan. As shown in Table 3, our model consistently outperform existing Tibetan instruction-tuned models across all tasks. In particular, Ours-SFT achieves the strongest overall performance, showing substantial improvements in reading comprehension and question answering, where other models fall significantly behind. Moreover, in both zh-bo and en-bo translation tasks, our models achieve the highest BLEU scores among all evaluated models, demonstrating strong multilingual alignment and effective cross-lingual generalization. It is worth noting that our comparison only includes open-source Tibetan-tailored models. Models such as Sun-Shine Huang et al. (2025), which are not publicly available, are not included in this evaluation.
@@ 7 Conclusion
To address the limited support for Tibetan in existing models, we construct a high-quality, large-scale Tibetan pre-training corpus. With the curated corpus, we perform continual pretraining to develop a multilingual large language model supporting Tibetan, Chinese, and English. Our model demonstrates superior performance compared to both general-purpose open-source models and Tibetan-tailored models of similar scale on our constructed Tibetan benchmark as well as traditional Tibetan NLP tasks.
@@ Appendix B Open-Source Data Collection Details
The open-source Tibetan corpus we collected primarily comes from two sources: (1) the Tibetan portions of multilingual open-source datasets, and (2) dedicated monolingual Tibetan datasets. These corpora cover a wide range of domains. However, we observe that a portion of the web-based data originates from different snapshots of Common Crawl, which may introduce redundancy. To address this, we appliy deduplication to remove duplicate content across time.
@@ Appendix C Web Data Crawling Details
We manually survey 45 Tibetan-language websites and collect web content from them. Most of these sites are news-oriented, and our primary goal is to obtain up-to-date Tibetan text with strong temporal relevance. During the web crawling process, we extract hyperlink elements from each page and determine whether the linked page belonged to the same root domain. If so, and the page has not been previously visited, we recursively crawl the subpage. Throughout the entire process, all crawling behavior strictly adhere to the rules defined in each site’s robots.txt file.
@@ Appendix D Quality Filter Details
To eliminate politically sensitive content from the collected data, we manually curated a blacklist of sensitive terms. Any document containing one or more terms from this list is filtered out.
@@ Appendix E Continual Pre-training Details
To improve the training efficiency of the the model, we adopt a packing strategy Raffel et al. (2020) to preprocess the data. This involves concatenating and truncating the data to ensure that each training sample is exactly 4096 tokens in length. Given the large volume of training data, we also apply a pre-tokenization strategy, where the tokenized and packed datasets are preprocessed and saved in advance to streamline the subsequent training process.
@@ Appendix F Long Context Pre-training Details
Data Distribution. To fully account for the distributional characteristics of the training data in the first stage and further enhance the model’s long context processing capabilities in Tibetan, we introduce longer mixed-language texts (Chinese, English, and Tibetan) as training corpora in the second training stage.We sample 0.56B tokens of Chinese data, 2.5B tokens of English data, and 1.5B tokens of Tibetan data (unused in the first stage) from long context sources in each respective language, resulting in a trilingual long context training set totaling 4.56B tokens.
@@ Appendix G Post Training Details
We extensively collect open-source instruction datasets in Chinese, English, and Tibetan. Upon analysis, we find that most existing Tibetan instruction datasets are generated by translating open-source Chinese and English instruction data using Google Translate. While this approach is simple and convenient, it often results in low-quality Tibetan instructions due to the current limitations of Google’s Tibetan translation capabilities. Additionally, such translations may introduce unintended biases. To address these issues, we collaborate with native Tibetan speakers to produce additional high-quality translation and summarization data. These samples are organized using instruction-style templates to further enhance the model’s instruction-following abilities in Tibetan, particularly in the areas of translation and summarization.